OpenStudy (anonymous):

Find g(x) if g(x) is the resulting function from moving f(x) = (x + 2) right 3 units and up 4 units.

OpenStudy (anonymous):

i got:g(x) = (x + 4) + 5

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

g(x) = x +2 (-3 this is to the right) +4 (this is up) = x+3

OpenStudy (haseeb96):

@loser why here -3 will come and will become this equation x+3 instead of x+9?

OpenStudy (anonymous):

wait so i was wrong? @loser66

OpenStudy (anonymous):

@cwrw238

OpenStudy (loser66):

that's the rule of shifting function. Keep in mine: to the right , you minus the amount of unit (not +)

OpenStudy (loser66):

@Haseeb96 is it clear?

OpenStudy (anonymous):

ohh so it should be g(x) = (x + 4) + 3

OpenStudy (loser66):

Your f(x) = x+2 now, move it 3 units to the right, you have f(x-3) = (x-3) +2= x-1, right? now, move it up 4 units, that means the whole thing +4 (x-1) +4 = x +3, that is g(x) which is the result when we move f(x) to the right, then up.

OpenStudy (loser66):

https://www.desmos.com/calculator/ubalnxxenq the red is original f(x) move 3 units to the right to get the blue one.

OpenStudy (loser66):

then up 4units to get the green one

OpenStudy (loser66):

https://www.desmos.com/calculator/es7iooangd

OpenStudy (anonymous):

im so confused

OpenStudy (loser66):

the first graph, I mark on x-intercept for you to see how the x-intercept move from the original one to a middle new one the second graph, I mark on y-intercept for you to see how it up is.

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

so it is +4

OpenStudy (loser66):

you see it??

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

i got the answer:g(x) = (x – 1) + 4

OpenStudy (anonymous):

g(x) = (x–1) + 4

OpenStudy (anonymous):

g(x) = (x minus 1) + 4

OpenStudy (loser66):

yes, and need simplifying. :)